OpenStudy (anonymous):

Question: -5^(n-7) -2=-3 Answer: -log5^(16+5)/7... How?

OpenStudy (jeyekomon):

-5^(n-7) -2=-3 5^(n-7) +2=3 5^(n-7) =1 log( 5^(n-7) ) = log(1) (n-7) * log(5) = log(1) (n-7) * log(5) = 0 n-7 = 0 n=7

OpenStudy (anonymous):

Thank you but that wasn't the answer listed on the HW. The answer was: (-log5)16+5 all divided by 7

OpenStudy (jeyekomon):

Well, then your answer has to be incorrect. I have no idea, how (-log5)16+5 / 7 would be the solution..

OpenStudy (anonymous):

I'll double check. Thank you though :)